# On the symmetry energy expansion and the peak value of the bulk viscosity

### v1

@@ IV.1 Peak value of bulk viscosity
Figure 1: The maximum value of the bulk viscosity $\zeta^{\text{max}}$ at $1n_{0}$, for a 1 kHz density oscillation. Left panel: $\zeta^{\text{max}}$ contour plot in the $JL$ plane. Overlaid are the Essick et al. constraints (their Fig. 1, 90% credible region) [72] and the three $L=L(J)$ relations from Eq. 8a. Right panel: $\zeta^{\text{max}}$ as a function of $L$, using each of the three $L=L(J)$ lines, with the Essick et al. and PREX-II [71] constraints overlaid.
@@
$\zeta^{\text{max}}=\dfrac{1}{2\omega}\left|\dfrac{A^{2}}{B}\right|.$ (21)
The value of the bulk viscosity at the peak is set by the EoS (and the hydrodynamics through $\omega$). With Eq. 20, we can calculate the maximum bulk viscosity as a function of density. The simplest expression is the maximum value of the bulk viscosity at $n_{B}=n_{0}$, which is
@@
The left panel of Fig. 1 displays a contour plot of the maximum bulk viscosity at $n_{0}$ in the $JL$ plane. A density oscillation with a linear frequency $f=1\text{ kHz}$ (which is a typical frequency expected in neutron star mergers [34]) is assumed. First, we see that, indeed, the bulk viscosity sharply dips to zero along the line $J=L$. As one departs from the line $J=L$, the value of $\zeta^{\text{max}}$ grows. In particular, the maximum bulk viscosity is strongly dependent on the value of $L$, which is interesting given the controversy surrounding the value of $L$ at present. Overlaid on the plot in pink is the 90% credible region for $J$ and $L$ from Essick et al. [72]. Also overlaid are the three correlation lines (Eq. 8a): green represents the median value and the gray and blue curves encompass 95.4% of the EoSs as discussed near Eq. 8a. To avoid confusion, I note that the results of Essick et al. and the three correlation lines from Tews et al. are unrelated: the former is determined from a Bayesian analysis of $\chi$EFT, astrophysical measurements, and PREX-II data, while the latter represents correlations discovered within a large set of physically acceptable Skyrme and RMF EoSs.
@@
On the right hand side of Fig. 1, I plot the bulk viscosity as a function of $L$, with $J$ as a function of $L$ taking the values given in the 3 lines in the left panel. Also displayed is the 90% credible region for $L$ from Essick et al. The left panel indicates that the Essick et al. results allow for the existence of a conformal point at $n_{B}=n_{0}$. However, once $J$ is marginalized over in the Essick et al. analysis, and once my $J$ values are “collapsed” onto $L$ values with the three correlations in Eq. 8a, the right panel indicates that a conformal point at $n_{B}=n_{0}$ is unlikely. Of course, there is much room for improvement in this comparison.
@@
Figure 2: The resonant-maximum value of the bulk viscosity $\zeta^{\text{max}}$ as a function of density, assuming a 1 kHz density oscillation. The different curves correspond to various choices of $L$. The other symmetry coefficients $J$, $K_{\text{sym}}$, and $Q_{\text{sym}}$ are chosen according to the correlations Eqs. 8a, 8b, and 8c using the median values.
However, Fig. 1 indicates that it is most likely that the resonant peak of bulk viscosity at saturation density is around $10^{27}$ or $10^{28}$ g/(cm s). Of course, if the central value of $L$ deduced from the PREX-II result is the correct one and $L$ is very large, than the peak value of bulk viscosity can be several times larger. Overall, the picture I present here is consistent with the results of Yang et al. [64], who found that small changes in $L$ result in large changes in the bulk viscosity.
@@
It is also interesting to look at how the peak bulk viscosity changes as density increases. Eqs. 20 and 21 indicate that $K_{\text{sym}}$ and $Q_{\text{sym}}$ appear in the $\zeta^{\text{max}}$ expression if one pushes away from $n_{B}=n_{0}$. In Fig. 2, I plot the maximum bulk viscosity as a function of density, for a few chosen values of $L$. The other symmetry coefficients are chosen according to Eqs. 8a, 8b, and 8c. It is important to remember that the bulk viscosity as a function of density does not look like Fig. 2. This figure instead shows, if one plots the bulk viscosity in the density-temperature plane, the height of the peak if one traces it from low density to high density. The actual behavior of $\zeta(n_{B},T_{0})$ at some fixed temperature $T_{0}$ depends on $\gamma$. Fig. 2 shows that, generally, the bulk viscosity is larger as density increases. The higher $L$ is, the faster this growth is. Note that this is different from the observation made in, for example, [101, 61], which found that the bulk viscosity increases with $L$. They concluded, correctly, that the beta-equilibration rate increases with the proton fraction, which increases with $L$, and this increases the bulk viscosity (for temperatures below the resonant peak temperature). In this paper, I am focusing only on the resonant-peak value of the bulk viscosity, which depends on $L$ (among other symmetry parameters) due to the behavior of the susceptibilities and does not involve the beta-equilibration rate directly. In any case, as density increases the bulk viscosity could exceed the values shown in Fig. 1. Calculations of the bulk viscosity for several RMF models in [35, 44, 49] confirm the tendency of the maximum bulk viscosity to rise with density.
@@ IV.2 Damping timescale for local density oscillations
A useful measure of the effect of bulk viscosity is the timescale for bulk viscosity to damp a small-amplitude density oscillation. The amplitude of the oscillation cancels out, which makes this quantity particularly useful. The energy density of an oscillation in nuclear matter is given by
@@
$\tau_{\text{diss}}^{\text{min}}=\dfrac{n_{B}\dfrac{\partial P}{\partial n_{B}}\bigg{|}_{x_{p}}}{\omega^{2}\zeta_{\text{max}}}=\dfrac{2}{\omega}\left(\dfrac{\dfrac{\partial P}{\partial n_{B}}\bigg{|}_{x_{p}}}{\dfrac{\partial P}{\partial n_{B}}\bigg{|}_{x_{p}}-\dfrac{\partial P}{\partial n_{B}}\bigg{|}_{\delta\mu}}\right).$ (26)
The incompressibility is the density derivative of the pressure (Eq. 10) and is evaluated in beta equilibrium, yielding
$\dfrac{\partial^{2}\varepsilon}{\partial n_{B}^{2}}\bigg{|}_{x_{p}}$ $=\dfrac{1}{n_{B}}\dfrac{\partial P}{\partial n_{B}}\bigg{|}_{x_{p}}=\dfrac{1}{9n_{0}}\bigg{\{}\left(K+K_{\text{sym}}+6L\right)\left(1+\dfrac{12J-24L-4K_{\text{sym}}}{K+K_{\text{sym}}+6L}x_{p}^{\text{eq.}}+...\right)$ (27)
@@
This second derivative is essentially the spring constant of dense matter. Notably, while for symmetric matter (e.g., nuclei), this spring constant would be just $K$, in beta-equilibrated dense matter, the relevant quantity becomes $K+K_{\text{sym}}+6L$, which can be significantly different from K (for high $L$, double or triple the size).
@@
×$\times\left\{1+4\left(\dfrac{3J+6L+K_{\text{sym}}+2K}{K+K_{\text{sym}}+6L}\right)x_{p}^{\text{eq.}}\right\}.$
For simplicity, I keep only the value of $\tau_{\text{diss}}^{\text{min}}$ at $n_{0}$, though one can calculate it to linear order in $(n_{B}-n_{0})/n_{0}$ without introducing higher order powers than $Q_{\text{sym}}$ in the symmetry energy.
In Fig. 3, I plot the minimum damping time in the $K_{\text{sym}}L$ plane, with $J$ chosen according to Eq. 8a and $K=250\text{ MeV}$. The key feature here is that the damping time is the inverse of the bulk viscosity. The bulk viscosity nearly vanishes (at $n_{0}$) when $J\approx L$ and thus the damping time goes to infinity. As $L$ increases, the maximum bulk viscosity grows dramatically and thus the damping time shrinks. Superimposed on this is the effect that matter with higher $L$ and $K_{\text{sym}}$ is stiffer, and therefore an oscillation of a certain amplitude contains more energy than that same oscillation would in softer matter and thus the dissipation time is longer for the same bulk viscosity. The overall message of this plot is that unless $L$ is quite close to $J$, rapid bulk-viscous damping can occur in neutron stars mergers if there are fluid elements that experience bulk viscosity at the resonant maximum (which itself is a key question for the relevance of bulk viscosity in neutron star mergers [37, 45, 3]). The Essick et al. constraints depicted in Fig. 3 indicate that the damping times could be just under 10 ms, but could also be as long as several hundred ms.
@@ IV.3 Peak rotational frequency of neutron stars and the r-mode instability window
Fig. 4 shows a schematic of this $T\Omega$ plane. Along the x-axis, $\Omega=0$ and the star in question is stable, as no r-modes are generated. Even for very small $\Omega$, as long as there is any viscosity at all, the rotating star is stable. However, as the rotation frequency grows larger, the gravitational-instability timescale drops dramatically and the star will be unstable. Clearly, there is some border between stable and unstable stars, and two contrasting possibilities (for purely $npe$-matter stars) are sketched in blue and green in Fig. 4. The existence of hyperonic, quark, or superfluid phases, or dark matter, inside the neutron star leads to more complicated features [104, 105, 27, 106, 107], but I do not consider these possibilities here.
@@
Since the Urca rate $\gamma$ depends relatively little on density, an isothermal neutron star at the right temperature could experience resonant bulk viscosity across its entire volume, with the value of the bulk viscosity being $\zeta(r)=\zeta^{\text{max}}(r)$ with $\zeta^{\text{max}}$ (c.f. Eq. 21). That is, at this one temperature, the bulk viscosity in the star is entirely a function of $A^{2}/B$ and $\omega_{r}$. A star at this temperature experiences more bulk-viscous dissipation than a star at any other temperature and thus this temperature supports the largest neutron star rotation rate $\Omega$ that is stable to r-modes. This idea is depicted in Fig. 4, where the two black dots correspond to the maximally-spinning neutron star discussed above, in the case where the maximum spin rate is less than (blue curve) or greater than (green curve) the Kepler frequency. I have not found any definitive statement in the literature about which case is realized in nature, though most papers assume the green curve scenario. My goal here is to use the results from previous sections to determine whether the maximum stable rotation rate is above or below the Kepler frequency, as a function of the symmetry energy parameters.
@@
Figure 5: Contour plot of the maximum stable rotation frequency of a $1.4M_{\odot}$ $npe$-matter neutron star, normalized by the Kepler frequency, as a function of symmetry energy parameters $L$ and $K_{\text{sym}}$. $J$ is set according to Eq. 8a. Only bulk-viscous damping is considered. The shaded parameter space corresponds to an r-mode instability window such as that shown by the blue curve in Fig. 4, while the unshaded parameter space has an instability window like that in the green curve in Fig. 4.
@@
In Fig. 5, I show a contour plot of the maximum spin frequency of a neutron star, normalized by the Kepler frequency, as a function of $L$ and $K_{\text{sym}}$. $J$ is chosen according to Eq. 8a. Small bulk viscosities correspond to small maximum stable rotation rates, and large bulk viscosities correspond to large maximum stable rotation rates, often above the Kepler frequency and thus not physically realizable. Fig. 5 indicates that there is a sizable pocket of symmetry energy parameter space that gives rise to a star with a maximum bulk viscosity that is still quite small, and thus only allows a star to spin stably at 100%, 50%, or even as low as 20% of its Kepler frequency (the blue region in this plot corresponds to the blue curve in Fig. 4). This is consistent with the results of Figs. 1 and 2 which indicate that very small resonant-peak values for bulk viscosity are still allowed by experiment.
@@
The temperature at which this local maximum in the maximum stable rotation rate occurs cannot be determined by the EoS, just as the temperature where the bulk viscosity reaches its resonant maximum could not. Calculations indicate in $npe$ matter, for a 1 kHz oscillation, the resonance occurs between temperatures of 2 and 6 MeV [35, 44]. Depending on the mass of the star, and its rotation rate, the (corotating) r-mode frequency $\omega_{r}$ would not be 1 kHz, but instead only a few hundred Hz. However, this only results in a slight ($\sim$ 1 MeV) decrease in the temperature of the resonance.
@@
At such high temperatures, while one can define an r-mode instability window within the paradigm discussed in this work (a rigidly rotating, isothermal neutron star made purely of $npe$ matter), complications come into the picture. A piece of $npe$ matter, cooling via modified Urca neutrino emission, cools to below 1 MeV in just a few seconds (see the simple model developed in [108]). The timescale for the growth of the r-mode due to gravitational radiation is as small as 10 s (Eq. 32 or see Fig. 4 of [27]). A star located just above one of the two black dots in Fig. 4 will, in all likelihood, cool before it feels the instability. In addition, neutron stars at temperatures this high are likely not isothermal, as the thermal equilibration of a neutron star core takes a decade or more [109, 110]. However, a non-isothermal star would just have a smaller fraction of its volume experiencing maximal bulk viscosity than the extreme case considered here. Finally, at temperatures above several MeV, neutrinos become trapped and locally equilibrated with the $npe$ matter [89, 111] and thus the EoS must include their contribution as well.

### v3

@@ IV.1 Peak value of bulk viscosity
Figure 1: The maximum value of the bulk viscosity $\zeta^{\text{max}}$ at $1n_{0}$, for a 1 kHz density oscillation. Left panel: $\zeta^{\text{max}}$ contour plot in the $JL$ plane. Overlaid are the Essick et al. constraints (their Fig. 1, 90% credible region) [74] and the three $L=L(J)$ relations from Eq. 8a. Right panel: $\zeta^{\text{max}}$ as a function of $L$, using each of the three $L=L(J)$ lines, with the Essick et al., PREX-II, and CREX constraints overlaid. For PREX-II, I show the original Reed et al. constraint [73] and the reanalysis by Koehn et al. [77]. For CREX, I show just the analysis by Koehn et al. [77]. The black dots indicate the median values.
@@
$\zeta^{\text{max}}\left(n_{B}\right)=\dfrac{1}{2\omega}\left|\dfrac{A^{2}}{B}\right|.$ (21)
The value of the bulk viscosity at the peak is set by the EoS (and the hydrodynamics through $\omega$). With Eq. 20, one can calculate the maximum bulk viscosity as a function of density. The simplest expression is the maximum value of the bulk viscosity at $n_{B}=n_{0}$, which is
@@
[…]
@@
On the right hand side of Fig. 1, I plot the bulk viscosity as a function of $L$, with $J$ as a function of $L$ taking the values given in the 3 Tews et al. lines in the left panel. Also displayed is the 90% credible region for $L$ from Essick et al. The left panel indicates that the Essick et al. results allow for the existence of a conformal point at $n_{B}=n_{0}$. However, once $J$ is marginalized over in the Essick et al. analysis, and once my $J$ values are “collapsed” onto $L$ values with the three correlations in Eq. 8a, the right panel indicates that a conformal point at $n_{B}=n_{0}$ is unlikely. Of course, more sophisticated analyses of the likelihood of a conformal point should be done.
The right panel of Fig. 1 also shows the predicted values of the symmetry energy slope $L$ coming from PREX-II and CREX. The PREX-II values, from Reed et al. and later from Koehn et al., predict a wide range for $L$, but the median values are quite high, around 100 MeV. In contrast, CREX predicts a much lower value of $L$. If PREX’s median values are correct, than the bulk viscosity at saturation is large, around $10^{29}$ g/(cm s), while if the true value of $L$ is around the CREX median value, a conformal point of the EoS at $n_{B}=n_{0}$ is possible and the bulk viscosity could be quite small. Thus, the tension between PREX and CREX has significant implications for the bulk viscosity at densities near saturation density. However, if the more conventional values of $L$ are correct (represented by the Essick et al. credible interval), then the bulk viscosity at saturation density is likely around $10^{27}$ or $10^{28}$ g/(cm s). Overall, the picture I present here is consistent with the results of Yang et al. [64], who found that small changes in $L$ result in large changes in the bulk viscosity.
@@
Figure 2: The resonant-maximum value of the bulk viscosity $\zeta^{\text{max}}$ as a function of density, assuming a 1 kHz density oscillation. The different curves correspond to various choices of $L$. The other symmetry coefficients $J$, $K_{\text{sym}}$, and $Q_{\text{sym}}$ are chosen according to the correlations Eqs. 8a, 8b, and 8c using the median values. Overlaid are the ranges of $\zeta^{\text{max}}$ predicted at $n_{0}$ from the values of $L$ found by Essick et al. [74], and the PREX-II [73, 77] and CREX [77] data.
It is also interesting to look at how the peak bulk viscosity changes as density increases. Eqs. 20 and 21 indicate that $K_{\text{sym}}$ and $Q_{\text{sym}}$ appear in the $\zeta^{\text{max}}$ expression if one pushes away from $n_{B}=n_{0}$. In Fig. 2, I plot the maximum bulk viscosity as a function of density, for a few chosen values of $L$. The other symmetry coefficients are chosen according to Eqs. 8a, 8b, and 8c. Also shown are the Essick et al., PREX-II, and CREX predictions for the bulk viscosity at $n_{B}=n_{0}$, given the range of $L$ values that they each predict.
@@
It is important to remember that the bulk viscosity as a function of density does not look like Fig. 2. This figure instead shows, if one plots the bulk viscosity in the density-temperature plane, the height of the peak if one traces it from low density to high density. The actual behavior of $\zeta(n_{B},T_{0})$ at some fixed temperature $T_{0}$ depends on $\gamma$. Fig. 2 shows that, generally, the bulk viscosity is larger as density increases. The higher $L$ is, the faster this growth is. Note that this is different from the observation made in, for example, [106, 61], which found that the bulk viscosity increases with $L$. These works concluded, correctly, that the beta-equilibration rate increases with the proton fraction, which increases with $L$, and this increases the bulk viscosity (for temperatures below the resonant peak temperature). In this paper, I am focusing only on the resonant-peak value of the bulk viscosity, which depends on $L$ (among other symmetry parameters) due to the behavior of the susceptibilities and does not involve the beta-equilibration rate directly.
In any case, as density increases the bulk viscosity likely (but not necessarily) exceeds the values shown in Fig. 1. Calculations of the bulk viscosity for several RMF models in [35, 44, 49] confirm the tendency of the maximum bulk viscosity to rise with density. Again, the value of $L$ is crucial in determining the behavior of the peak bulk viscosity value as a function of density.
@@ IV.2 Damping timescale for local density oscillations
A useful measure of the effect of bulk viscosity is the timescale for bulk viscosity to damp a small-amplitude density oscillation. The amplitude of the oscillation cancels out, which makes this quantity particularly useful. The energy density of an oscillation in nuclear matter is given by [^5] One could presumably choose, for the purpose of estimating the oscillation energy, to keep $\delta\mu$ fixed instead of $x_{p}$. The difference between the adiabatic and equilibrium incompressibilities is small compared to either one of them (c.f. Eq.26).
@@
$\tau_{\text{diss}}^{\text{min}}=\dfrac{n_{B}\dfrac{\partial P}{\partial n_{B}}\bigg|_{x_{p}}}{\omega^{2}\zeta_{\text{max}}}=\dfrac{2}{\omega}\left(\dfrac{\dfrac{\partial P}{\partial n_{B}}\bigg|_{x_{p}}}{\dfrac{\partial P}{\partial n_{B}}\bigg|_{x_{p}}-\dfrac{\partial P}{\partial n_{B}}\bigg|_{\delta\mu}}\right),$ (26)
where Eq. 21 and 18 were used. The incompressibility is the density derivative of the pressure (Eq. 10) and is evaluated in beta equilibrium, yielding
[…]
@@
This second derivative is essentially the spring constant of dense matter. Notably, while for symmetric matter without leptons (e.g., nuclei), this spring constant would be just $K$, in beta-equilibrated dense matter, the relevant quantity becomes $K+K_{\text{sym}}+6L$, which can be significantly different from K (for high $L$, double or triple the size).
@@
$\times\left\{1+4\left(\dfrac{3J+6L+K_{\text{sym}}+2K}{K+K_{\text{sym}}+6L}\right)x_{p}^{\text{eq.}}\right\}.$
For simplicity, I display only the value of $\tau_{\text{diss}}^{\text{min}}$ at $n_{0}$, though one can calculate it to linear order in $(n_{B}-n_{0})/n_{0}$ without introducing higher order powers than $Q_{\text{sym}}$ in the symmetry energy.
In Fig. 3, I plot the minimum damping time in the $K_{\text{sym}}L$ plane, with $J$ chosen according to Eq. 8a and $K=250\text{ MeV}$. The key feature here is that the damping time is the inverse of the bulk viscosity. The bulk viscosity vanishes (at $n_{0}$) when $J=L$ and thus the damping time goes to infinity. As $L$ increases, the maximum bulk viscosity grows dramatically and thus the damping time shrinks. Superimposed on this is the effect that matter with higher $L$ and $K_{\text{sym}}$ is stiffer, and therefore an oscillation of a certain amplitude contains more energy than that same oscillation would in softer matter and thus the dissipation time is longer for the same bulk viscosity. The overall message of this plot is that unless $L$ is quite close to $J$, rapid bulk-viscous damping can occur in neutron stars mergers if there are fluid elements that experience bulk viscosity at the resonant maximum (which itself is a key question for the relevance of bulk viscosity in neutron star mergers [37, 45, 3]). The Essick et al. constraints depicted in Fig. 3 indicate that the damping times could be just under 10 ms, but could also be as long as several hundred ms. Of course, if the PREX-II median values of $L$ are correct, then damping could be as quick as a few ms, whereas if the CREX median value is correct, then the damping time would be quite long.
@@ IV.3 Peak rotational frequency of neutron stars and the r-mode instability window
Fig. 4 shows a schematic of this $T\Omega$ plane. Along the x-axis, $\Omega=0$ and the star in question is stable, as no r-modes are generated. Even for very small $\Omega$, as long as there is any viscosity at all, the rotating star is stable. However, as the rotation frequency grows larger, the gravitational-instability timescale drops dramatically and the star will be unstable. Clearly, there is some border between stable and unstable stars, and two contrasting possibilities (for purely npenpe-matter stars) are sketched in blue and green in Fig. 4. The existence of hyperonic, quark, or superfluid phases, or dark matter, inside the neutron star leads to more complicated features [109, 110, 27, 111, 112], but I do not consider these possibilities here.
@@
Since the Urca rate $\gamma$ depends relatively little on density, an isothermal neutron star at the right temperature could experience resonant bulk viscosity across its entire volume, with the value of the bulk viscosity being $\zeta(r)=\zeta^{\text{max}}(r)$ with $\zeta^{\text{max}}$ given by Eq. 21. That is, at this one temperature, the bulk viscosity in the star is entirely a function of $A^{2}/B$ and $\omega_{r}$. A star at this temperature experiences more bulk-viscous dissipation than a star at any other temperature and thus this temperature supports the largest neutron star rotation rate $\Omega$ that is stable to r-modes. This idea is depicted in Fig. 4, where the two black dots correspond to the maximally-spinning neutron star discussed above, in the case where the maximum spin rate is less than (blue curve) or greater than (green curve) the Kepler frequency. I have not found any definitive statement in the literature about which case is realized in nature, though most papers seem to assume the green curve scenario. My goal here is to use the results from previous sections to determine whether the maximum stable rotation rate is above or below the Kepler frequency, as a function of the symmetry energy parameters.
@@
Figure 5: Contour plot of the maximum stable rotation frequency of a $1.4M_{\odot}$ npenpe-matter neutron star, normalized by the Kepler frequency, as a function of symmetry energy parameters $L$ and $K_{\text{sym}}$. $J$ is set according to Eq. 8a. Only bulk-viscous damping is considered. The shaded parameter space corresponds to an r-mode instability window such as that shown by the blue curve in Fig. 4, while the unshaded parameter space has an instability window like that in the green curve in Fig. 4.
@@
In Fig. 5, I show a contour plot of the maximum spin frequency of a neutron star, normalized by the Kepler frequency, as a function of $L$ and $K_{\text{sym}}$. The symmetry coefficient $J$ is chosen according to Eq. 8a. Small bulk viscosities correspond to small maximum stable rotation rates, and large bulk viscosities correspond to large maximum stable rotation rates, often above the Kepler frequency and thus not physically realizable. Fig. 5 indicates that there is a sizable pocket of symmetry energy parameter space that gives rise to a star with a maximum bulk viscosity that is still quite small, and thus only allows a star to spin stably at 100%, 50%, or even as low as 20% of its Kepler frequency (the blue region in this plot corresponds to the blue curve in Fig. 4). This is consistent with the results of Figs. 1 and 2 which indicate that very small resonant-peak values for bulk viscosity are still allowed by experiment.
@@
The temperature at which this local maximum in the maximum stable rotation rate occurs cannot be determined by the EoS, just as the temperature where the bulk viscosity reaches its resonant maximum could not. Calculations indicate in npenpe matter, for a 1 kHz oscillation, the resonance occurs between temperatures of 2 and 6 MeV [35, 44]. Depending on the mass of the star, and its rotation rate, the (corotating) r-mode frequency $\omega_{r}$ would not be 1 kHz, but instead only a few hundred Hz. However, this only results in a slight ($\sim$ 1 MeV) decrease in the temperature of the resonance.
@@
At such high temperatures, while one can define an r-mode instability window within the paradigm discussed in this work (a rigidly rotating, isothermal neutron star made purely of npenpe matter), complications come into the picture. A piece of npenpe matter, cooling via modified Urca neutrino emission, cools to below 1 MeV in just a few seconds (see the simple model developed in [113]). The timescale for the growth of the r-mode due to gravitational radiation is as small as 10 s (Eq. 32 or see Fig. 4 of [27]). A star located just above one of the two black dots in Fig. 4 will, in all likelihood, cool before it feels the instability. In addition, neutron stars at temperatures this high are likely not isothermal, as the thermal equilibration of a neutron star core takes a decade or more [114, 115]. However, a non-isothermal star would just have a smaller fraction of its volume experiencing maximal bulk viscosity than the extreme case considered here. Finally, at temperatures above several MeV, neutrinos become trapped and locally equilibrated with the npenpe matter [86, 116] and thus the EoS must include their contribution as well.